Like tan 3# = 3tan# - tan3#/1 -3tan2#


then how to find tan 4#, tan 5#, tan 6#...........
so
we donot hve to memorize it till last


:consider # as thitha angle’


 



Grade:12th Pass

1 Answers

eaishwary bajpai
12 years ago

solve tan4# as taking

4# = 2(2#)

 

tan(5#) = tan (2# + 3#)

 

tan(6#) = tan2(3#)